OpenStudy (anonymous):

A salesperson finds that her sales average 45 cases per store when she visits 5 stores a week. Each time she visits an additional store per week, the average sales per store decreases by 1 case. How many stores should she visit each week if she wants to maximize her sales?

OpenStudy (anonymous):

4?

OpenStudy (turingtest):

9=45/5 C= cases s=number of stores n=additional number of stores the condition given can be written as \[9-n={C \over s+n}\]that's just me thinking out loud... don't have the answer yet.

OpenStudy (anonymous):

take the derivadtive

OpenStudy (anonymous):

of your function and find the maximum and minimum and you will get the answer

OpenStudy (anonymous):

good job turing test thumbs up for you

OpenStudy (anonymous):

We're on polynomial functions in my class... Wouldn't we look for h and k using these formulas: h=-B/2A and k=C-B^2/4A?

OpenStudy (anonymous):

Idk...

OpenStudy (anonymous):

let's derive a formula

OpenStudy (turingtest):

great, then I can't use calculus... either way gotta get C in terms of n though I think

OpenStudy (turingtest):

sorry another typo \[9-n={45\over5+n}\]

OpenStudy (anonymous):

is not 45 should be 5

OpenStudy (anonymous):

you have the the average

OpenStudy (anonymous):

no you are right

OpenStudy (turingtest):

C/s=9=45/5 since C/s goes down by the number of additional stores\[{45\over5}-n={45\over5+n}\to 9-n={45\over9+n}\]so I'm pretty sure the formula is right, now let's get it quadratic looking...

OpenStudy (anonymous):

now just how did you 9+n

OpenStudy (anonymous):

there should be n+1

OpenStudy (anonymous):

or n-1 i meant if n represents additional number of store

OpenStudy (anonymous):

where did you guys even get this formula?

OpenStudy (turingtest):

sorry I keep typing 5 when I mean 9\[9-n={45\over5+n}\]is the formula

OpenStudy (anonymous):

not the #s, just the formula where u plugged in the numbers

OpenStudy (anonymous):

listen n represents number of additional store which we have in the question but there is a decrease of 1

OpenStudy (turingtest):

9 is the average sales for each of the stores (C/s=45/5) in our initial condition, so in general rate of sales=cases sold/number of stores R=C/s ever time the # of stores goes up by one the rate goes down, so r-n=C/(s+n) that is where my formula came from.

OpenStudy (turingtest):

now lets get it quadratic-looking: (9-n)(5+n)=45+4n-n^2=45 4n-n^2=0 which as a maximum at n=2, so the total number of stores to visit is 5+2=7

OpenStudy (anonymous):

Ok... We didn't talk about this in class but it makes sense.

OpenStudy (turingtest):

I think you can get the max from one of the formulas you used h=-b/2a=-4/2(-1)=2

OpenStudy (anonymous):

wait shouldnt it be -14n instead of 4n?

OpenStudy (anonymous):

-9n + -5n = -14n

OpenStudy (turingtest):

(9-n)(5+n)=45+9n-5n-n^2=45+4n-n^2

OpenStudy (anonymous):

o nvm u changed one of the signs to a positive

OpenStudy (anonymous):

thats not the answer

OpenStudy (anonymous):

neither 2 , 4, or 7

OpenStudy (turingtest):

hmm...

OpenStudy (anonymous):

Can you solve this prob using: f(x)=ax^2+bx+c f(x)=a(x-h)+k h=-b/2a k=c-(b^2/4a)

OpenStudy (turingtest):

well not if I set it up wrong, lol

OpenStudy (anonymous):

the 2nd one should be a(x-h)^2+k (my bad)

OpenStudy (anonymous):

Cause that's what we've been doing so far in class... Idk why the hw would be different.

OpenStudy (turingtest):

You could use that formula on my equation and get n=2 so total # of stores=7 but you say that is wrong, so it must be set up wrong.

OpenStudy (anonymous):

1 and 5 arent the answers either... just tried those out

OpenStudy (turingtest):

not sure how: say she visits 6 stores. One more store means the average goes down to 8, so total sales is 56 say she visits 7 stores. Now the number of stores is 7 and the rate is 7, so total is 49. if she visits 8 stores then the rate is 6 so the total sales is 56. hmmm, it looks like 6 stores and 8 stores are both the maximum?

OpenStudy (turingtest):

that can't be... what happened?

OpenStudy (anonymous):

6 * 8 = 48

OpenStudy (turingtest):

oh dear, right so we get 7 again, which is what I originally said...

OpenStudy (turingtest):

How is it not seven? hmm... I must not be seeing something drastic :(

OpenStudy (turingtest):

Oh no! I see I misread 45 is the AVERAGE, not 9

OpenStudy (turingtest):

\[45-n={225 \over5+n}\]that should do it...

OpenStudy (anonymous):

So far these are not the answers: 1-5 and 7

OpenStudy (turingtest):

no solve this one 225+40n-n^2=225 n=20 for maximum total stores=25 Please tell me that's right

OpenStudy (anonymous):

yes it is

OpenStudy (turingtest):

thank god, do you see how I got it and what mistake I made?

OpenStudy (anonymous):

where did you get 225?

OpenStudy (turingtest):

total sales/stores=rate of sales per store C/s=R C/5=45 C=225 now the formula I had all along will work:\[R-n={C \over s+n}\]\[45-n={225\over5+n}\]All I did was misread that 45 was the total, not the average.

OpenStudy (anonymous):

Ok makes sense... If only my teacher taught us what we need to know for the homework. -_-

OpenStudy (anonymous):

wait howd u get 25... im working it out on paper and i get to -n^2+40n=0

OpenStudy (turingtest):

\[(45-n)(5+n)=225+45n-5n+n^2=225+40n-n^2\]

OpenStudy (anonymous):

yeah i got 225+40n-n^2=225 and i subtracted 225 so i just got 40n-n^2

OpenStudy (turingtest):

\[225+40n-n^2=225\]is what we have after multiplying both sides by 5+n which reduces to \[40n-n^2=0\]now use your formula for h=-b/2a

OpenStudy (anonymous):

o

OpenStudy (turingtest):

we find n=20 so total stores=5+n=25

OpenStudy (anonymous):

ok i see. thanks.